# Request-Only Optimization for Recommendation Systems

### v1

@@ Abstract.
Deep Learning Recommendation Models (DLRMs) represent one of the largest machine learning applications on the planet. Industry-scale DLRMs are trained with petabytes of recommendation data to serve billions of users every day. To utilize the rich user signals in the long user history, DLRMs have been scaled up to unprecedented complexity, up to trillions of floating-point operations (TFLOPs) per example. This scale, coupled with the huge amount of training data, necessitates new storage and training algorithms to efficiently improve the quality of these complex recommendation systems.
In this paper, we present a Request-Only Optimizations (ROO) training and modeling paradigm. ROO simultaneously improves the storage and training efficiency as well as the model quality of recommendation systems. We holistically approach this challenge through co-designing data (i.e., request-only data), infrastructure (i.e., request-only based data processing pipeline), and model architecture (i.e., request-only neural architectures). Our ROO training and modeling paradigm treats a user request as a unit of the training data. Compared with the established practice of treating a user impression as a unit, our new design achieves native feature de-duplication in data logging, consequently saving data storage. Second, by de-duplicating computations and communications across multiple impressions in a request, this new paradigm enables highly scaled-up neural network architectures to better capture user interest signals, such as Generative Recommenders (GRs) and other request-only friendly architectures.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Neural networksccs: Information systems Recommender systems
@@ 1. Introduction
The critical issue emerges at the impression-level event-feature joiner (Figure 1.b), where extensive feature duplication occurs. For each user request that generates multiple impressions, identical user features are redundantly stored across all impression-level training samples. Figure 2 reveals that industrial recommendation products typically generate 4-7 impression samples per request. This redundancy cascades throughout the system - consuming storage, network bandwidth, and GPU computation resources.
@@
While Request-Only Optimization (ROO) has been recently adopted for model inference (You et al., 2024), extending this concept to training presents significant challenges. Existing ROO implementations often act as wrappers around impression-level data at inference time, increasing system complexity. Potential training-time optimization is complicated by the widespread use of impression-based data formats throughout client and server infrastructures, and the incremental arrival of impression-based events during the feedback phase.
@@
Building on this foundation, we develop ROO training methodologies and ROO-friendly model architectures that capitalize on this data structure. By processing user-side information at the request level rather than the impression level, we reduce the computational workload from $B_{NRO}$ to $B_{RO}$ examples (where $B_{NRO}$ represents the number of impression items in a batch and $B_{RO}$ represents the number of requests). With an average of 4-7 impressions per request ($B_{NRO}/B_{RO}=4-7$) and user-side features dominating the feature space (Figure 2), our approach yields theoretical speedups of 300-600%.
@@
  • A new ROO training paradigm that maximizes DLRM system efficiency through computation redundancy elimination and at the same time, unifying the data format used across training and inference tech stacks.
@@ 2. End-to-End ROO Design
In this work, we take a first-principled approach and address the fundamental problem at the root level of data for DLRM training and inference. The end-to-end Request-Only Optimization training paradigm benefits from a cohesive co-design with data, system, and model architecture. The well-designed data format eliminates the feature duplication at the source and enables us to hatch a systematical implementation throughout storage, GPU resource usage, training, and inference. The synergy between ROO and advanced modeling techniques, such as Generative Recommenders (Zhai et al., 2024), strikes a great balance between model complexity and efficiency.
@@ 2.1.2. ROO Data Activity-to-Serving Latency
In recommendation products, activity-to-serving (ATS) latency is used to measure personalization models’ freshness, and activity-to-training latency is a critical part of it. In theory, adopting request-level joining could potentially increase activity-to-training latency because it needs to wait for multiple items within the request join window. In production, request-level joiners use either fixed-time policy or dynamic training trigger to close the join window before the fixed-time window closes. Through measurements, we establish that the average time gap between the earliest user-item interaction event arriving at the request join window buffer and the last item of the request is within 16 minutes. In practice, ROO training samples’ data landing latency is about the half the fixed-time join window and satisfies model ATS requirements.
2.1.3. ROO Expansion for Backward Compatibility
Despite its significant storage efficiency advantage, request-level training sample schema is incompatible with the traditional DLRMs trained with impression-level schema. In a typical industrial environment, hundreds of models across multiple ranking stages collectively serve billions of users using a product, and not all models have to adopt the ROO training paradigm due to return-on-investment (ROI) considerations. We present a practical way to minimize the migration cost for these models and facilitate the adoption of ROO data schema. We introduce ROO expansion data adapter in the feature preprocessing layer. For each mini-batch, the data adapter expands a ROO sample into multiple in-memory impression samples, with user-side features duplicated to the in-memory data frames. Trainer side rebatching is leveraged to keep the consistent training batch size. This equips ROO data schema with backward compatibility and eliminates unnecessary model rewrite efforts. The expansion functionality simplifies the migration process; major products fully migrated all retrieval, early stage ranking and late stage ranking models to ROO data within a 9-month time frame.
@@ 2.2. Unifying Model Training and Inference Request-Level Optimization
Figure 3. Compared to impression training, ROO training deduplicates RO-side embedding lookup, embedding all2all communication, RO-only model arch’s computation FLOPs from $B_{NRO}$ batch size to $B_{RO}$ batch size. In the example, user and item embedding tables are sharded across GPU-0 and GPU-1 accordingly; $B_{NRO}$ is 4 and $B_{RO}$ is 1.
@@
In a trainer mini-batch, the user-side RO and item-side NRO feature tensors have different batch sizes. The RO tensor batch size $B_{RO}$ is the number of ROO training samples, while the NRO tensor batch size $B_{NRO}$ is the sum of impression items across all request level samples in the trainer mini-batch. This allows trainers to leverage TorchRec’s variable-length batch size feature (Ivchenko et al., 2022) for model parallel between RO and NRO computation. The system efficiency gains primarily come from the user-side RO computation deduplication. User-side RO feature embedding lookup and associated all2all communication overhead are reduced to exactly once for each request-level sample in the mini-batch. Compared to impression level training, as illustrated in Figure 3, the computation and communication complexity is reduced from $O(B_{NRO})$ to $O(B_{RO})$. In our extensive experiments in DLRM systems (Section 4), the RO model arch reduced overall FLOPs significantly as a result of the ROO-aware training paradigm.
@@ 3.3. ROO-based Autoregressive Modeling
ROO enables scalable autoregressive modeling, with a commonly used variant, generative recommenders (GRs) (Zhai et al., 2024; Zhang et al., 2025; Huang et al., 2025; Borisyuk et al., 2025; Han et al., 2025) as follows. Denote a sequence of $n$ past impressions in user history $RO_{0},RO_{1},$ $\ldots,RO_{n-1}(RO_{i}\in\mathbb{X})$ ordered chronologically, and a sequence of $m$ target items in the request $NRO_{0},NRO_{1},\ldots,NRO_{m-1}$ ($NRO_{i}\in\mathbb{X}$), where $\mathbb{X}$ denotes the set of all items served in the product. With each item associated with an action feature $a$ and various contextual features including but not limited to surface type, timestamp, denoted as $c$, sequential modeling in standard retrieval and ranking models can be formulated as shown in Figure 5(b).
@@ 4.1. End-to-End System Efficiency Savings
ROO training data captures all impressions of the same request in one single request-level training sample and hereby eliminates duplications at the fundamental data level. Through data and system co-design, we eliminated computational redundancy throughout training data pipeline, intra-datacenter network transfer, feature preprocessing, GPU training, and model inference. As a result, ROO training achieved multi-fold efficiency benefits.
ROO data allows logging more training samples and avoids downsampling with the same storage capacity. As shown in Figure 6, we observed 43% to 150% increases in training sample volumes across a few recommendation products.
Figure 6. Training sample volume increase across different recommendation products, eliminating data downsampling.
The adoption of ROO training optimization yields substantial benefits, including significant reductions in GPU training costs and notable increases in training throughputs. This breakthrough enables the exploration of previously cost-prohibitive modeling technologies, such as UserArch and GRs/HSTU, which were formerly hindered by computational constraints. To demonstrate the efficacy of ROO training, we conducted experiments on multiple models across various recommendation stages without modifying the model architecture. The results, presented in Figure 7, show an increase in training throughput of 48% to 570% in retrieval models, 125% to 266% in early stage ranking models, and 32% to 100% in late stage ranking models, achieved through our proposed ROO optimization technique.
Figure 7. Training throughput increase (in percentage) on multiple models across the recommendation stages.
@@ 4.2. Model Quality Improvements
We next conduct a range of experiments in the retrieval, early-stage ranking (ESR), and late-stage ranking (LSR) models from three billion-user scale recommendation products, hereinafter referred to as product-A, -B, and -C. These experiments validate the practicality of the ROO paradigm and the related model architecture innovations in production environments. We also provide detailed ablation studies and online A/B test results to showcase the impact from ROO in production.
@@ 4.3. Retrieval and Early-Stage Ranking Models
Model FLOPs per example C-Task Recall@100 E-Task Recall@100 Topline Consumption
@@ 4.3.1. Retrieval Models
Without the ROO data format, model scaling can contribute to improved offline results at the expense of 6.8x more computation costs per training example. With ROO, the same improvements in offline recall can be attained with no increase in training cost per item example as we now effectively amortize computational cost across multiple impressions. We collected online A/B test metrics were collected from a large-scale experiment on product-B. We observed a 4.77% increase in the main online consumption metric after launching the HSTU architecture on ROO.
Table 4. Normalized entropy (NE) improvement in offline experiments of ranking models from multiple product surfaces and online A/B test metrics improvements after the models are deployed to production. The significant level of each metric across different models are marked in separate rows. In the LSR model on product-A, UserArch and HSTU are tested and deployed sequentially so that their metrics are listed incrementally. E.g. the two modules brings 0.41% and 0.53% NE improvement (reduction) on E-task respectively. On product-B, two LSR models, 1 and 2, were tested.
@@
LSR-1 on product-B UserArch+HSTU -0.86% -0.78% +0.07% +5.56% +7.0%
LSR-2 on product-B UserArch+HSTU -0.60% -0.64% +0.04% +2.51% +6.9%
@@ 4.3.2. Early-stage Ranking (ESR) Models
With the ROO paradigm, we further integrated HSTU into the user tower of the ESR models. As shown in Table 4, HSTU improved the recommendation quality as measured both offline NE and post launch topline impact. Specifically, this design leads to significant boost of the main consumption metric by 2.43% and the main topline metric by 0.04% to product-B, where 0.24% and 0.03% are significant, respectively.
@@ 4.4. Late-Stage Ranking (LSR) Models
LSR on product-A. In this experiment, we conducted an offline ablation analysis on a late-stage ranking model for product-A. We incorporated two ROO architectures into the model: UserArch and HSTU. On this model, we deployed UserArch to production first, followed by HSTU, which was tested and rolled out subsequently. From both offline and online experiments, HSTU can add notable incremental improvement to the model on top of UserArch. Table 4 summarizes the offline evaluation NE improvement and post-launch online metrics lift. Note that a 0.1% NE improvement on both engagement and consumption tasks is considered substantial for this particular model, and 0.03% lift in the topline metric, 0.63% lift in engagement metric, and 0.06% lift in consumption metric are deemed significant for this product.
LSR on product-B. We conducted a comprehensive evaluation of UserArch and HSTU on two distinct models, LSR-1 and LSR-2, within the recommendation service for product-B. In this product, 0.03% lift in the topline metric, 0.63% lift in engagement metric, and 0.24% lift in consumption metric is considered as significant. Our offline and online experiments revealed that these ROO-related technologies yielded notable improvements in normalized entropy (NE) and post-launch online metrics in both models as shown in Table 4.
LSR on product-C. We tested HSTU on a LSR model for product-C, yielding significant improvements in NE. Specifically, we observed 0.81% and 0.79% NE reductions on engagement and consumption tasks, respectively, which are larger than the 0.1% significant level. Following the successful deployment of this model to production, we measured substantial gains in online metrics, including a 0.25% lift in topline metric and a 2.85% increase in engagement metric as shown in Table 4. Normally, 0.03% lift in topline metric, 0.19% in engagement metric and 0.04% in consumption metric are considered significant for this product.

### v3

@@ Abstract.
Recommendation systems represent one of the largest machine learning applications on the planet – industry-scale recommendation models are trained with petabytes of data and serve billions of users every day. To utilize the rich user signals in the long user history, these models have been scaled up to unprecedented complexity, up to trillions of floating-point operations (TFLOPs) per example. This scale, coupled with the huge amount of training data, necessitates new storage and training algorithms to efficiently improve the quality of these complex recommendation systems.
In this paper, we present a Request-Only Optimizations (ROO) training and modeling paradigm. ROO simultaneously improves the storage and training efficiency as well as the model quality of recommendation systems. We holistically approach this challenge through co-designing data (i.e., request-only data), infrastructure (i.e., request-only based data processing pipeline), and model architectures. Our ROO training and modeling paradigm treats a user request as a unit of the training data. Compared with the established practice of treating a user impression as a unit, this new design directly removes redundant features in data logging, saving data storage. Second, by de-duplicating computations and communications across multiple impressions in a request, this new paradigm enables highly scaled-up neural network architectures to better capture user interest signals, such as Generative Recommenders (GRs) and other request-only friendly architectures.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Neural networksccs: Information systems Recommender systems22footnotetext: Work done at Meta.
@@ 1. Introduction
The critical issue emerges at the impression-level event-feature joiner (Figure 1.b), where extensive feature duplication occurs. For each user request that generates multiple impressions, identical user features are redundantly stored across all impression-level training samples. Figure 2 reveals that industrial recommendation products typically generate 4-7 impression samples per request. This redundancy cascades throughout the system – consuming storage, network bandwidth, and GPU computation resources.
@@
While request-level optimizations has been recently applied to model inference (You et al., 2024) to exploit the single-user, multiple-candidates characteristic of recommendation model inference, extending this concept to training presents significant challenges. Existing implementations often act as wrappers around impression-level data at inference time, increasing system complexity. Potential training-time optimization is complicated by the widespread use of impression-based data formats throughout client and server infrastructures, and the incremental arrival of impression-based events during the feedback phase.
@@
Building on this foundation, we develop ROO training methodologies and ROO-friendly model architectures that capitalize on this data structure. By processing user-side information at the request level rather than the impression level, we reduce the computational workload from $B_{NRO}$ to $B_{RO}$ examples (where $B_{NRO}$ represents the number of impression items in a batch and $B_{RO}$ represents the number of requests). With an average of 4-7 (Figure 2) impressions per request ($B_{NRO}/B_{RO}$) and user-side features dominating the feature space, our approach yields theoretical speedups of 300-600%.
@@
  • A new ROO training paradigm that maximizes recommendation system efficiency through computation redundancy elimination and at the same time, unifying the data format used across training and inference tech stacks.
@@ 2. End-to-End ROO Design
In this work, we take a first-principled approach and address the fundamental problem at the root level of data for DLRM training and inference. The end-to-end Request-Only Optimization training paradigm benefits from a cohesive co-design with data, system, and model architecture. The data format eliminates the feature duplication at the source and enables us to hatch a systematical implementation throughout storage, GPU resource usage, training, and inference. The synergy between ROO and advanced modeling techniques, such as HSTU (Zhai et al., 2024), strikes a great balance between model complexity and efficiency.
@@ 2.1.2. ROO Data Activity-to-Serving Latency
In recommendation products, activity-to-serving (ATS) latency is used to measure personalization models’ freshness, and activity-to-training latency is a critical part of it. In theory, adopting request-level joining could potentially increase activity-to-training latency because it needs to wait for multiple items within the request join window. In production, request-level joiners use either fixed-time policy or dynamic training trigger to close the join window before the fixed-time window closes. Through measurements, we establish that the average time gap between the earliest user-item interaction event arriving at the request join window buffer and the last item of the request is within 16 minutes. In practice, ROO training samples’ data landing latency is about the half the fixed-time join window.
2.1.3. Request Level Join Data Quality
Changing training data joining from impression-level to request-level could theoretically lead to training data distribution changes which might have an impact on the recommendation model quality. However, our data quality analysis shows user training sample and feature coverage parity with impression-level training data. Table 3 shows that distributions of important labels such as conversion and video view duration have negligible mismatch.
Label Mismatch Rate
Conversion Mismatch Product-A 0.03%
Conversion Mismatch Product-B 0.16%
Conversion Mismatch Product-C 0.023%
Video View Duration Mismatch Product-A 0.035%
Video View Duration Mismatch Product-B 1.07%
Video View Duration Mismatch Product-C 0.01%
Table 3. Request level join data quality as compared to impression level data
@@ 2.2. Unifying Model Training and Inference Request-Level Optimization
Figure 3. Compared to impression training, ROO training deduplicates RO-side embedding lookup, embedding all-to-all communication, RO-only model arch’s computation FLOPs from $B_{NRO}$ batch size to $B_{RO}$ batch size. In the example, user and item embedding tables are sharded across GPU-0 and GPU-1 accordingly; $B_{NRO}$ is 4 and $B_{RO}$ is 1.
@@
In a trainer mini-batch, the user-side RO and item-side NRO feature tensors have different batch sizes. The RO tensor batch size $B_{RO}$ is the number of ROO training samples, while the NRO tensor batch size $B_{NRO}$ is the sum of impression items across all request level samples in the trainer mini-batch. This difference in batch sizes can be effectively managed by leveraging TorchRec’s variable-length batch sharding, allowing the same embedding table to be shared between RO and NRO features.
The system efficiency gains primarily come from the user-side RO computation deduplication. User-side RO feature embedding lookup and associated all-to-all communication overhead are reduced to exactly once for each request-level sample in the mini-batch. Compared to impression level training, as illustrated in Figure 3, the computation and communication complexity is reduced from $O(B_{NRO})$ to $O(B_{RO})$. In our extensive experiments in DLRM systems (Section 4), the RO model arch reduced overall FLOPs significantly as a result of the ROO-aware training paradigm.
@@ 3.3. ROO-based Autoregressive Modeling
ROO enables scalable autoregressive modeling, with a commonly used variant, generative recommenders (GRs) (Zhai et al., 2024; Zhang et al., 2025; Huang et al., 2025; Borisyuk et al., 2025; Han et al., 2025; Khrylchenko et al., 2025) as follows. Denote a sequence of $n$ past impressions in user history $RO_{0},RO_{1},$ $\ldots,RO_{n-1}(RO_{i}\in\mathbb{X})$ ordered chronologically, their associated user actions (e.g., like or skip), and a sequence of $m$ target items in the request $NRO_{0},NRO_{1},\ldots,NRO_{m-1}$ ($NRO_{i}\in\mathbb{X}$), where $\mathbb{X}$ denotes the set of all items served in the product. With each item associated with an action feature $a$ and various contextual features including but not limited to surface type, timestamp, denoted as $c$, sequential modeling in standard retrieval and ranking models can be formulated as shown in Figure 5(b).
@@ 4.1. End-to-End System Efficiency Savings
ROO training data captures all impressions of the same request in one single request-level training sample and hereby eliminates duplications at the fundamental data level. Through data and system co-design, we eliminated computational redundancy throughout training data pipeline, intra-datacenter network transfer, feature preprocessing, GPU training, and model inference. As a result, ROO training achieved multi-fold efficiency benefits. First we benchmarked the efficiency savings from ROO through a range of experiments in the retrieval, early-stage ranking (ESR), and late-stage ranking (LSR) models from three billion-user scale recommendation products, hereinafter referred to as product-A, -B, -C, and -D.
ROO data allows logging more training samples and avoids downsampling with the same storage space. As shown in Table 4, we observed 43% to 150% increases in training sample volumes across a few recommendation products. The storage space saving depends heavily on the proportion of user (request) side features and the number of training examples within a request. Heterogeneous user behaviors across different recommendation products leads to different training example distributions. In the extreme case, if a request contains only one training example, ROO data effectively reverts to impression level data, and no de-duplication is possible.
Training samples increase
Product-A 43%
Product-B 150%
Product-C 100%
Table 4. Training sample volume increase across different recommendation products.
ROO training significantly boosts training throughput, enabling the exploration of previously compute-prohibitive modeling technologies like UserArch and GRs/HSTU. We benchmarked ROO and impression-level training throughput across multiple models and recommendation stages. In all experiments, the training data originated from the same industrial recommendation source, and the neural network architectures remained consistent. The only controlled variations were the ROO data, preprocessing, and training pipeline.
As shown in Table 5, our proposed ROO techniques achieved substantial increases in training throughput: 220% to 570% in retrieval models, 125% to 266% in early-stage ranking models, and 32% to 100% in late-stage ranking models. In practice, these throughput improvements are also influenced by the percentage of compute and memory that can be de-duplicated within individual models. Two-tower models with a heavy user tower, such as retrieval and ESR models., benefit more from ROO than the LSR models.
In recommendation product-D surface, we compared a RecD-trained ESR model with ROO-trained model where ROO training showed 50% more training throughput.
Model Training throughput increase
LSR Product-A 50%
ESR Product-A 125%
LSR Product-B 33%
ESR Product-B 266%
Retrieval Product-B 570%
LSR Product-C 32%
ESR Product-C 150%
Retrieval Product-C 220%
Table 5. Training throughput increase (in percentage) on multiple models across the recommendation stages.
@@ 4.2. Model Quality Improvements
We next conduct experiments to validate the practicality of the ROO paradigm and the related model architecture innovations in production environments. We also provide detailed ablation studies and online A/B test results to showcase the impact from ROO in production.
@@ 4.3. Retrieval and Early-Stage Ranking Models
Model
FLOPs per
example
C-Task
Recall@100
E-Task
Recall@100
Topline Consumption
@@ 4.3.1. Retrieval Models
Without the ROO data format, model scaling can contribute to improved offline results at the expense of 6.8x more computation costs per training example. With ROO, the same improvements in offline recall can be attained with no increase in training cost per item example as we now effectively amortize computational cost across multiple impressions. We collected online A/B test metrics from a large-scale experiment on product-B. We observed a 4.77% increase in the main online consumption metric after launching the HSTU architecture on ROO.
Table 7. Normalized entropy (NE) improvement in offline experiments of ranking models from multiple product surfaces and online A/B test metrics improvements after the models are deployed to production. The significant level of each metric across different models are marked in separate rows. In the LSR model on product-A, UserArch and HSTU are tested and deployed sequentially so that their metrics are listed incrementally. E.g. the two modules brings 0.41% and 0.53% NE improvement (reduction) on E-task respectively.
@@
LSR on product-B UserArch+HSTU -0.86% -0.78% +0.07% +5.56% +7.0%
@@
LSR on product-D UserArch+HSTU -0.60% -0.64% +0.04% +2.51% +6.9%
Significant delta on product-D 0.1% 0.1% 0.03% 0.24% 0.63%
@@ 4.3.2. Early-stage Ranking (ESR) Models
With the ROO paradigm, we further integrated HSTU into the user tower of the ESR models. As shown in Table 7, HSTU improved the recommendation quality as measured by both offline NE and post launch topline impact. Specifically, this design leads to significant boost of the main consumption metric by 2.43% and the main topline metric by 0.04% to product-B, where 0.24% and 0.03% are significant, respectively.
@@ 4.4. Late-Stage Ranking (LSR) Models
LSR on product-A. In this experiment, we conducted an offline ablation analysis on a late-stage ranking model for product-A. We incorporated two ROO architectures into the model: UserArch and HSTU. On this model, we deployed UserArch to production first, followed by HSTU, which was tested and rolled out subsequently. From both offline and online experiments, HSTU added notable incremental improvement to the model on top of UserArch. Table 7 summarizes the offline evaluation NE improvement and post-launch online metrics lift. Note that a 0.1% NE improvement on both engagement and consumption tasks is considered substantial for this particular model, and 0.03% lift in the topline metric, 0.63% lift in engagement metric, and 0.06% lift in consumption metric are deemed significant for this product.
LSR on product-B. We conducted a evaluation of UserArch and HSTU on an LSR model within the recommendation service for product-B. In this product, 0.03% lift in the topline metric, 0.63% lift in engagement metric, and 0.24% lift in consumption metric is considered as significant. Our offline and online experiments revealed that these ROO-related technologies yielded notable improvements in NE and post-launch online metrics in both models as shown in Table 7.
LSR on product-C. We tested HSTU on a LSR model for product-C, yielding significant improvements in NE. Specifically, we observed 0.81% and 0.79% NE reductions on engagement and consumption tasks, respectively, which are larger than the 0.1% significant level. In online A/B test, we observed substantial gains in online metrics, including a 0.25% lift in the primary topline metric and a 2.85% increase in engagement metric as shown in Table 7. Normally, 0.03% lift in topline metric, 0.19% in engagement metric and 0.04% in consumption metric are considered significant for this product.
LSR on product-D. On product-D, we tested UserArch and HSTU on an LSR model. We have observed distinct improvements in both offline and online experiments exceeding the significant level as shown in Table 7. Particularly, in offline experiment, 0.60% and 0.64% NE reductions on engagement and consumption tasks were observed. In online A/B test, the topline, consumption, and engagement metrics were lifted by 0.04%, 2.51%, and 6.9% respectively.